# Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation

### v1

@@ 6.3.3 Detailed results from the loss ablation study
In the main paper, we provide ablation of loss components and their combinations i.e., the mean accuracy across all the tasks for ViT-B/16 and ViT-B/32 on the benchmarks corruption datasets. Here, we provide additional task-wise accuracy in Table 12. The addition of loss components $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ help CLIP in adapting its feature space to a specific domain/corruption.
@@ 6.3.4 Post-adaptation results on source test sets
In this crucial experiment, we challenge our BAT-CLIP, and evaluate its zero-shot generalization performance back on the source test set, to check the preservation of pre-trained CLIP Knowledge. After the adaptation of CLIP on each corruption task, we report the adapted model’s zero-shot performance on its corresponding source test set. We report results for CIFAR-10C and CIFAR-100C in Tables 10 and 11, using ViT-B/16 and ViT-B/32 backbones. For all of the results, we use the prompt template “a photo of a ¡CLS¿.”. As an example, for CIFAR-10C, upon adaptation of CLIP to Gaussian noise following our approach, we report the adapted model’s zero-shot accuracy on its source test set - CIFAR10 test set.
@@
On the whole, we conclude that since the adaptation for a task happens over multiple test batches, the zero-shot performance back on the source data largely depends on the distribution of the image corruption. Overall, ViT-B/16 visual backbones preserve larger amounts of CLIP pre-trained knowledge. This proves the effectiveness of our method BAT-CLIP, on average.
@@
In this section, we provide illustrations of task-wise t-SNE [43] plots for CIFAR-10C and CIFAR-100C and compare them against zero-shot ViT-B/16. The results are in Figures 10, 11, 12, and 13. Across all corruptions/tasks, BAT-CLIP learns strong discriminative visual features with a strong image-text alignment and class-level separation. ImageNet-C has 1000 classes, so, we do not provide t-SNE plots to avoid complications. However, the analysis and results carry forward.
@@
Figure 10: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
@@
Figure 12: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.

### v2

@@ 6.3.3 Detailed results from the loss ablation study
In the main paper, we provide ablation of loss components and their combinations i.e., the mean accuracy across all the tasks for ViT-B/16 on the benchmark corruption datasets. Here, we provide additional task-wise accuracy in Table 6. The addition of loss components $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ help CLIP in adapting its feature space to a specific domain/corruption.
@@
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
90.1
Ours
ViT-B/32
88.3
Ours
Table 7: Zero-shot performance on CIFAR10 (source) after adaptation of BATCLIP on a task.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
66.6
Ours
ViT-B/32
62.3
Ours
Table 8: Zero-shot performance on CIFAR100 (source) after adaptation of BATCLIP on a task.
6.3.4 Effect of different prompt templates
In Table , we show results with “relevant” prompt templates to show the independence of such prompt selection, at test-time. As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates. Though TPT [45] fine-tunes a pre-trained prompt on each image, and VTE [10] uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time usage.
In all of our prior experiments, we use a generic prompt template “a photo of a $<$CLS$>$.” for all of the datasets and methods. Here, we replace this with “relevant” prompt templates to show the independence of such a prompt selection, at test-time, and report the results in Table 9. As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates. Though TPT [45] fine-tunes a pre-trained prompt on each test image, and VTE [10] uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time deployment.
Prompt Template CIFAR-10C CIFAR-100C ImageNet-C
a low contrast photo of a $<$CLS$>$.
68.53 (+7.81) 37.09 (+4.97) 27.31 (+3.71)
a blurry photo of a $<$CLS$>$.
68.84 (+10.96) 36.80 (+5.33) 26.92 (+3.52)
a photo of a big $<$CLS$>$.
67.49 (+10.10) 35.79 (+4.87) 25.64 (+3.29)
6.3.5 Impact of bimodal adaptation
To show the effectiveness of bimodal adaptation, we ablate $\mathcal{L}_{pm}$, which is responsible for updating the text encoder $f_{txt}$. We report the results in Table 10. We see a drop in accuracy when $f_{txt}$ is “frozen” i.e., when $\mathcal{L}_{pm}$ isn’t used, necessitating the need for bimodal adaptation of CLIP encoders.
Dataset $f_{vis}$ update $f_{txt}$ update Ours
CIFAR-10C 72.58
73.85
CIFAR-100C 41.00
42.09
ImageNet-C 29.88
30.72
@@ 6.3.6 Post-adaptation results on source test sets
In this crucial experiment, we challenge our BATCLIP, and evaluate its zero-shot generalization performance back on the source test set, to check the preservation of pre-trained CLIP Knowledge. After the adaptation of CLIP on each corruption task, we report the adapted model’s zero-shot performance on its corresponding source test set. We report results for CIFAR-10C and CIFAR-100C in Tables 7 and 8, using ViT-B/16 and ViT-B/32 backbones. For all of the results, we use the prompt template “a photo of a $<$CLS$>$.”. As an example, for CIFAR-10C, upon adaptation of CLIP to Gaussian noise following our approach, we report the adapted model’s zero-shot accuracy on its source test set - CIFAR10 test set.
@@
On the whole, we conclude that since the adaptation for a task happens over multiple test batches, the zero-shot performance back on the source data largely depends on the distribution of the image corruption. Overall, ViT-B/16 visual backbones preserve larger amounts of CLIP pre-trained knowledge. This proves the effectiveness of our method BATCLIP, on average.
@@
In this section, we provide illustrations of task-wise t-SNE [49] plots for CIFAR-10C and CIFAR-100C and compare them against zero-shot ViT-B/16. The results are in Figures 11, 12, 13, and 14. Across all corruptions/tasks, BATCLIP learns strong discriminative visual features with a strong image-text alignment and class-level separation. ImageNet-C has 1000 classes, so, we do not provide t-SNE plots to avoid complications. However, the analysis and results carry forward.
@@
Figure 11: BATCLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
@@
Figure 13: BATCLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.